OpenStudy (anonymous):

How do you integrate this using substitution ? ∫x^3(5+x^2)^1/2dx

OpenStudy (mimi_x3):

\[\int x^{3}\sqrt{5+x^2}dx\]

OpenStudy (mimi_x3):

let u = \(x^2\) =>du/dx = \(2x\)

OpenStudy (lgbasallote):

hmm

OpenStudy (anonymous):

\[∫ 1/(x²+4)^3/2 dx = 1/4 \sin \quad u + C\] \[∫ 1/(x²+4)^3/2 dx = x/(4√(x²+4)) + C \]

OpenStudy (mimi_x3):

\[\int\limits x^{3} \sqrt{5+u} *\frac{du}{2x} => \int\limits x*x^{2} \sqrt{5+u} *\frac{du}{2x} => \int\limits x*u\sqrt{5+u} *\frac{du}{2x} \] \[ =>\frac{1}{2} \int\limits u\sqrt{5+u} \]

OpenStudy (mimi_x3):

Are you able to do it now?

OpenStudy (mimi_x3):

@lgbasallote: What's wrong?

OpenStudy (anonymous):

nothing wrong with that

OpenStudy (anonymous):

@Mimi_x3 Why did you write u in place of x^2 ? :/

OpenStudy (anonymous):

nvm got it .thanks alot :)

OpenStudy (mimi_x3):

np

OpenStudy (anonymous):

sorry one more question @Mimi_x3 , why did you set u = x^2 . Isnt it always the inside function so in this case u = 5+x^2 ?

OpenStudy (callisto):

I would do it in this way... \[\int x^3 \sqrt{5+x^2} dx\]Let u = 5+x^2 du = 2x dx ; x^2 = u-5 \[\int x^3 \sqrt{5+x^2} dx\]\[=\frac{1}{2}\int x^2 \sqrt{5+x^2} (2xdx)\]\[=\frac{1}{2}\int (u-5) \sqrt{u} du\]\[=\frac{1}{2}\int u^{\frac{3}{2}}-5u^{\frac{3}{2}} du\]\[=...\]

OpenStudy (callisto):

*for the last step before ..., it is \[=\frac{1}{2} \int u^{\frac{3}{2}}-5u^{\frac{1}{2}}du\]

OpenStudy (anonymous):

thankyou :D

OpenStudy (mimi_x3):

yeah but you have to change the variable.. any way would work